OpenStudy (anonymous):

Solve the following equation: (p+4)(p-3) = 8 p=

OpenStudy (nikato):

First FOIL (p+4)(p-3)

OpenStudy (anonymous):

p^2-3p+4p-12=8

OpenStudy (nikato):

Ok. Now combine like terms

OpenStudy (anonymous):

(p-4)(p+5)=0

OpenStudy (nikato):

Ok. Correct. And can you find p now?

OpenStudy (anonymous):

p=-5,4

TheSmartOne (thesmartone):

p+4=8 p=4 p-3-8 p=11

OpenStudy (nikato):

Yes

TheSmartOne (thesmartone):

So p=11 or p=4

OpenStudy (nikato):

p=-5 or 4 is correct

OpenStudy (anonymous):

thanks!

OpenStudy (nikato):

No problem